### v2

# LLaSE-G1: Incentivizing Generalization Capability for LLaMA-based Speech Enhancement

Abstract
Recent advancements in language models (LMs) have demonstrated strong capabilities in semantic understanding and contextual modeling, which have flourished in generative speech enhancement (SE). However, many LM-based SE approaches primarily focus on semantic information, often neglecting the critical role of acoustic information, which leads to acoustic inconsistency after enhancement and limited generalization across diverse SE tasks. In this paper, we introduce LLaSE-G1, a LLaMA-based language model that incentivizes generalization capabilities for speech enhancement. LLaSE-G1 offers the following key contributions: First, to mitigate acoustic inconsistency, LLaSE-G1 employs continuous representations from WavLM as input and predicts speech tokens from X-Codec2, maximizing acoustic preservation. Second, to promote generalization capability, LLaSE-G1 introduces dual-channel inputs and outputs, unifying multiple SE tasks without requiring task-specific IDs. Third, LLaSE-G1 outperforms prior task-specific discriminative and generative SE models, demonstrating scaling effects at test time and emerging capabilities for unseen SE tasks. Additionally, we release our code and models to support further research in this area  [^1] LLaSE-G1 Codes and Demos.
LLaSE-G1: Incentivizing Generalization Capability for LLaMA-based Speech Enhancement
Boyi Kang1 thanks: Equal contribution., Xinfa Zhu1 11footnotemark: 1, Zihan Zhang1 11footnotemark: 1, Zhen Ye2, Mingshuai Liu1, Ziqian Wang1, Yike Zhu1, Guobin Ma1, Jun Chen3, Longshuai Xiao3, Chao Weng3, Wei Xue2, Lei Xie1 1 Northwestern Polytechnical University 2 The Hong Kong University of Science and Technology 3 Huawei Technologies Co., Ltd. {beaukang, xfzhu, zhzhang, lxie}@mail.nwpu.edu.cn
1 Introduction
In recent years, large language models (LLMs) have made significant strides in natural language processing (NLP) OpenAI (2024), computer vision (CV) Tschannen et al. (2024); Chang et al. (2022), and speech processing Wang et al. (2023a); Zhang et al. (2023b), driving the rapid development of artificial intelligence technologies. In the NLP domain, LLMs have redefined text generation benchmarks through innovative pre-training and post-training paradigms, particularly excelling in few-shot and zero-shot learning scenarios. The impact of LLMs extends beyond unimodal textual processing. In CV research, integrating LLMs with visual models has sparked the rise of multimodal learning frameworks, facilitating more efficient processing of tasks such as image comprehension and generation. Similarly, the convergence of modalities is evident in the speech domain, where LLMs have enhanced the naturalness and accuracy of speech interaction systems. These advancements not only highlight the power of LLMs within individual domains but also underscore their potential for multimodal tasks.
Task Type Distortion Reference Signal
NS Noise, Reverb -
PLC Noise, Packet Loss Lossy Label
TSE
Noise, Reverb,
Interfering Speech
Enrolled Speech
AEC Noise, Reverb, Echo Echo Speech
SS
Noise, Reverb,
Interfering Speech
Table 1: Subtasks Definition in Speech Enhancement
As a fundamental task in the field of speech processing, speech enhancement (SE) aims to remove interference from noisy speech and separate and reconstruct clean target speech. Depending on the differences between the interfering speech and the target speech, sub-tasks can be defined as Noise Suppression (NS), Packet Loss Concealment (PLC), Target Speaker Extraction (TSE), Acoustic Echo Cancellation (AEC), Speech Separation (SS), and others, as detailed in Table 1. Neural SE models can generally be categorized into two types: discriminative Zhao et al. (2024a, c) and generative Wang et al. (2024). Deep learning-based discriminative SE models learn a mapping between degraded speech and the corresponding clean speech target. In contrast, generative SE models employ language models or diffusion models to learn the data distribution of the target speech. Notable recent models, including SELM Wang et al. (2024), TSELM Tang et al. (2024), and GenSE Yao et al. (2025), leverage semantic understanding and contextual modeling capabilities, achieving competitive performance in speech enhancement tasks. While traditional discriminative SE models require carefully designed architectures and task-specific loss functions, generative SE models offer a more flexible framework, enabling better scalability across different SE tasks.
Despite surpassing traditional discriminative models in speech quality, generative SE models still face challenges in acoustic preservation and task generalization. Many generative SE models rely on discrete speech tokens—typically extracted from speech codecs—as inputs to facilitate language modeling. However, as speech is inherently a continuous signal, using discrete tokens, especially semantic tokens, inevitably results in information loss Yao et al. (2025), leading to acoustic inconsistencies after enhancement, such as changes in speaker timbre and intonation. Moreover, most generative models are focused on a single task, such as noise suppression, which limits their generalization across different SE tasks. Since SE tasks differ in their input, output, and underlying functions, it remains an open question whether LMs can serve as versatile, multi-task SE models.
In this paper, we argue that, with appropriate design, a single language model can be a powerful and versatile SE model. To this end, we propose LLaSE-G1, a LLaMA-based language model that incentivizes generalization capabilities across various SE tasks. The architecture of LLaSE-G1 is simple yet effective, consisting of a WavLM Chen et al. (2022) encoder for feature extraction, a LLaMA-based language model for token prediction, and an X-codec2 Ye et al. (2025) decoder for waveform reconstruction. Specifically, to address the acoustic inconsistency caused by the information loss inherent in discrete tokens, we replace the discrete token inputs with continuous representations extracted from the WavLM encoder and predict speech tokens obtained from X-codec2. The WavLM encoder provides sufficient speech details, and X-codec2 integrates semantic and acoustic features into speech tokens, thus maximizing acoustic preservation. Additionally, to incentivize generalization, LLaSE-G1 utilizes dual-channel inputs and outputs, unifying the degraded speech and optional reference signals and constraining all tasks under a cross-entropy loss function. Through extensive experiments, LLaSE-G1 demonstrates superior performance on NS, PLC, TSE, and AEC benchmarks. Furthermore, LLaSE-G1 exhibits emergent capabilities for previously unseen SE tasks, such as SS, and shows scaling effects at test time, where performance improves with increased compute.
In summary, our paper makes several key contributions:
  • We propose LLaSE-G1, a LLaMA-based language model that incentivizes generalization capability for speech enhancement.
  • We effectively address the acoustic inconsistency by leveraging both continuous and discrete representations, and we design dual-channel inputs and outputs, which unify various SE tasks without the need for task IDs. Notably, AEC, PLC, and SS tasks are being introduced to generative models for the first time.
  • LLaSE-G1 outperforms existing models on several SE benchmarks and demonstrates scaling effects during test time and emergent capabilities for unseen SE tasks. We release the codes and checkpoints as open-source.
2 Related Work
Speech enhancement refers to the technology of recovering high-quality target speech from degraded speech, which includes multiple subtasks Wang and Chen (2018); Liu et al. (2022b). Traditional speech enhancement, which relies on statistical analysis and signal processing, often struggles with generalization in dynamic scenarios. With the development of deep learning, data-driven speech enhancement has become the mainstream approach and can be divided into two categories: discriminative SE and generative SE Lemercier et al. (2023). Discriminative SE models learn a mapping between degraded speech and the corresponding clean speech targets, including methods such as time-frequency (T-F) masking Williamson and Wang (2017) and time-domain approaches Luo and Mesgarani (2018). In contrast, generative models reconstruct the clean speech by learning the data distribution of the target speech, such as diffusion-based generative models Zhang et al. (2023a); Richter et al. (2023). Recently, researchers have also begun to explore the use of LMs to improve generative speech enhancement Wang et al. (2024); Yao et al. (2025).
2.1 Discriminative Speech Enhancement
Traditionally, speech enhancement encompasses tasks such as NS, PLC, TSE, AEC, and SS, with NS also requiring dereverberation. For different tasks, discriminative models often have different architectures. In NS tasks, most models are based on the convolutional encoder-decoder (CED) architecture. FRCRN Zhao et al. (2024a) adds a frequency recurrent network to the CED architecture, achieving excellent performance. In PLC tasks, Generative Adversarial Networks (GANs) are commonly used. BS-PLCNet Zhang et al. (2024) uses multitask learning and multi discriminators, winning the latest PLC Challenge Diener et al. (2024). In TSE tasks, the speaker embedding paradigm is widely adopted. TSE approaches usually use speaker verification models Desplanques et al. (2020); Wang et al. (2023b) to extract embeddings from enrollment audio and integrate into noise suppression networks. This approach has been successful in the personalized tracks of the Deep Noise Suppression challenges, as demonstrated by the TEA-PSE series models Ju et al. (2023, 2022). For AEC tasks, an important issue is how to deal with the delay estimation and alignment between reference signals and microphone signals. DeepVQE Indenbom et al. (2023a) utilizes attention-based delay estimation, employing fully neural networks to solve echo cancellation problems. For SS tasks, common models such as TF-GridNet Wang et al. (2023c) and Mossformer2 Zhao et al. (2024b) can only handle a fixed number of speakers. SepTDA Lee et al. (2024) introduces a transformer decoder-based attractor, capable of handling a dynamic number of speakers, but still requires specifying the maximum number of speakers.
While these discriminative models have achieved excellent performance across various tasks, their generalization ability is limited by the availability of training data and model parameters Welker et al. (2022). This can lead to performance degradation in unseen scenarios. Additionally, these models may introduce undesired speech distortion and phonetic inaccuracies to enhanced speech Wang et al. (2020).
2.2 Generative Speech Enhancement
Early generative SE primarily relied on GANs and VAEs Pascual et al. (2017); Fang et al. (2021). Although these approaches offered new perspectives, they still did not surpass the performance of discriminative models. In recent years, diffusion-based generative models have been applied to speech enhancement. CDiffusion Lu et al. (2022) defines the conditional diffusion process by incorporating noisy data into the diffusion process. DiffSep Scheibler et al. (2023) designs stochastic differential equations (SDE) Song et al. (2021). By solving the corresponding reverse-time SDE, it is possible to recover individual sources from the mixture. Despite diffusion models achieving superior speech quality over discriminative models in noise suppression (NS) and source separation (SS), these tasks were previously independent of each other, requiring separate training of different models, and proving difficult to generalize to other tasks.
Recently, researchers have begun to explore the use of a joint framework, leveraging the capabilities of generative models to integrate multiple enhancement tasks into a single model. Nemo Ku et al. (2024) and SpeechFlow Liu et al. (2024) pre-trained on large-scale datasets and can be adapted to downstream tasks such as NS and TSE through fine-tuning. AnyEnhance Zhang et al. (2025) achieves both NS and TSE without the need for fine-tuning. It introduces a prompt guidance mechanism, enabling in-context learning capabilities.
With the rise of LMs in their ability to handle multiple tasks, LMs have also been introduced into speech enhancement. SELM Wang et al. (2024) employs a WavLM-based k-means tokenizer and predicts clean tokens from noisy tokens, marking the first introduction of LMs into the NS domain. MaskSR Li et al. (2024) uses a mask generation model to simultaneously handle noise, reverberation, clipping, and bandwidth limitation. However, existing unified enhancement models have not considered the echo cancellation task, which requires reference audio input and the need to address delay estimation and alignment issues. We suggest that by leveraging the powerful modeling capabilities of LMs, it is possible to develop a general speech enhancement model that unifies NS, PLC, TSE, AEC and SS.
3 LLaSE-G1
3.1 Overall Architecture
LLaSE-G1 is designed to incentivize generalization across various SE tasks with a single LLaMA-based LM Grattafiori et al. (2024). As shown in Figure 1, compared to previous specialist models such as FRCRN Zhao et al. (2024a), TEA-PSE 3.0 Ju et al. (2023), Align-ULCNet Shetu et al. (2024b), mossformer2 Zhao et al. (2024c) and BS-PLCNet Zhang et al. (2024), LLaSE-G1 greatly simplifies the model structure, keeping three main components: (1) a WavLM encoder, (2) a LLaMA-based LM and (3) an X-codec2 decoder. Specifically, the WavLM encoder extracts continuous speech features from degraded speech. The LLaMA-based LM takes speech features as input and predicts discrete speech tokens extracted by X-codec2 in an autoregressive manner. Finally, the X-codec2 decoder reconstructs enhanced speech from predicted speech tokens.
Formally, let: 1. $\texttt{Vertorize}(X)=\{x_{1},\dots,x_{N}\}$ be the WavLM encoder, which converts input degraded speech $X$ into $N$ speech features. 2. $\texttt{Vertorize}(P)=\{p_{1},\dots,p_{T}\}$ be the WavLM encoder, which converts optional reference speech $X$ into $T$ speech features. 3. $\texttt{Tokenize}(Y)=\{y_{1},\dots,y_{M}\}$ be the X-codec2 encoder, which converts an enhanced speech $Y$ into $M$ speech tokens. 4. $\texttt{Detokenize}(\{y_{1},\dots,y_{M}\})=\hat{Y}$ be the X-codec2 decoder, which reconstructs the waveform $\hat{Y}$ from its token representations. As the downsampling rate of WavLM is the same as that of X-codec2, $N$ is equal to $M$. Given a dataset $\mathcal{D}=\{(X_{i},P_{i},Y_{i})\}_{i=1}^{S}$, where $X_{i}$ is the degraded speech, $Y_{i}$ is the enhanced speech and $P_{i}$ is the reference speech or empty if unavailable, we represent each pair $(X_{i},P_{i},Y_{i})$ as a sequence $(x_{1},\dots,x_{N},p_{1},\dots,p_{T},y_{1},\dots,y_{M})$. Since the $X_{i}$ and $P_{i}$ are always given as input during training and inference, we pad $X_{i}$ and $P_{i}$ to the same length and the LM $\theta$ focuses on learning the conditional probability:
$P(x_{1},\ldots,x_{N},p_{1},\ldots,p_{T},y_{1},\ldots,y_{M};\theta)$ (1)
$=$ $\prod_{j=1}^{M}P(y_{j}|x_{1}\odot p_{1},\ldots,x_{j}\odot p_{j};\theta),$
where $\odot$ is the concatenation between $x$ and $p$ in the channel axis.
Figure 1: Overall Architecture of LLaSE-G1. LLaSE-G1 simplifies model architecture to support various SE tasks.
3.2 Maximizing Acoustic Preservation
As highlighted by WavChat Ji et al. (2024), speech representations can be broadly categorized into two types: continuous and discrete representations. Continuous representations, typically extracted from self-supervised learning (SSL) models like HuBERT Hsu et al. (2021) and WavLM Chen et al. (2022), are considered lossless carriers for speech, capturing intricate speech details. In contrast, discrete representations are derived from speech codecs such as Encodec Défossez et al. (2022) and DAC Kumar et al. (2023), which, while facilitating language modeling, are lossy due to the information loss during quantization. To address this, LLaSE-G1 adopts continuous representations as input and predicts discrete representations, aiming to maximize acoustic preservation throughout the enhancement process.
For continuous speech representations, we utilize WavLM as the extractor. WavLM is an SSL model that combines a convolutional feature encoder with a transformer encoder. Pre-trained on large-scale speech data, it excels across various speech-processing tasks. Previous research Baas et al. (2023); Zhu et al. (2023) has shown that features extracted from the 6th layer of WavLM contain sufficient acoustic information for high-fidelity speech reconstruction. Therefore, we leverage the features from this layer as the input representations for the language model.
For discrete speech representations, we use X-Codec2 as the extractor. X-Codec2 is a recently developed speech codec that integrates semantic and acoustic features into a unified codebook, ensuring a 1D causal dependency. This design reflects the inherent left-to-right temporal structure of audio signals, while also preserving more acoustic information compared to traditional 1D semantic tokens. Consequently, we adopt the speech tokens extracted by X-Codec2 as the modeling target for the language model.
3.3 Unifying Various SE tasks
Although different speech enhancement tasks are applied in various scenarios, they share underlying commonalities, such as the need to determine which components should be removed from the noisy speech. To this end, LLaSE-G1 employs a dual-channel input and output framework that unifies several SE tasks within a single language model (LM). These tasks include NS, TSE, PLC, AEC, and SS.
Systematically, NS, PLC, and SS require only the degraded speech as input, while TSE and AEC need both the degraded speech and an additional reference speech. In contrast to previous PLC models, we do not use the lossy labels that indicate missing speech frames, thereby simplifying the data requirements.
To unify the input representations, we introduce a dual-channel input: one channel for the degraded speech and the other for the optional reference speech. These representations are padded to the same length and concatenated along the channel dimension. Notably, if the reference speech is unavailable, we set the second channel to zero.
While NS, PLC, AEC, and TSE tasks output a single enhanced speech, we note that AEC requires the removal of information related to the reference speech, while TSE necessitates the preservation of reference speech information. To address this, we introduce a dual-channel output with two linear projection heads to unify the output representations. The first channel $c_{1}$ predicts tokens related to reference speech and the second channel $c_{2}$ predicts tokens irrelevant to reference speech. With these designs, for tasks including NS, AEC, and PLC, we employ a single-supervision strategy $\mathcal{L}_{\text{S}}$ through the cross-entropy loss between $c_{0}$ and the tokens $t_{0}$ extracted from the clean signal:
$\mathcal{L}_{\text{S}}=-\frac{1}{N}\sum_{k=1}^{N}t_{0}^{(k)}\log\left(c_{0}^{(k)}\right)$ (2)
For the TSE task, we implement a dual-supervision strategy $\mathcal{L}_{\text{D}}$ with separate constraints for both outputs. The first output $c_{0}$ handles interfering speaker, while the second output $c_{1}$ is dedicated to target speaker extraction. The $\mathcal{L}_{\text{D}}$ is formulated as:
$\mathcal{L}_{\text{D}}=-\frac{1}{N}\sum_{k=1}^{N}[t_{0}^{(k)}\log\left(c_{0}^{(k)}\right)+t_{1}^{(k)}\log\left(c_{1}^{(k)}\right)]$ (3)
Importantly, in LLaSE-G1, SS is treated as an unseen task throughout the entire training process.
4 Experiments and Results
4.1 Experimental Setup
Datasets. For the training data, we use the Librispeech, HiFiTTS, and DNS Challenge datasets Reddy et al. (2020); Dubey et al. (2023), along with internal datasets as original clean speech, totalling approximately 5000 hours. Room impulse responses (RIRs) are sourced from the DNS Challenge datasets. The noise data contains nearly 1000 hours, sourced from DEMAND, ESC-50, DNS Challenge, AEC Challenge Cutler et al. (2023), and internal datasets.
Data augmentation. We utilized dynamic data augmentation during training. For the NS task, the clean audio and noise are randomly selected and mixed with a signal-to-noise Ratio (SNR) ranging from [-5,20] dB. Both clean and noisy signals have a 50% probability of adding reverberation. In the PLC task, we use a two-state first-order Markov chain to describe the packet loss status of the current frame and the next frame. The transition and hold probabilities for Markov states are selected between 0.05 and 0.95. We directly generate a binary mask sequence and apply it to the clean speech. For the AEC task, we randomly select a real echo signal and its corresponding reference signal from the far-end single talk in the AEC Challenge dataset. The signal-to-echo ratio (SER) ranges from -15 dB to 15 dB. Noise is added with a 20% chance, and the SNR is between -5 dB and 20 dB. For the Target Speech Enhancement (TSE) task, we select a clean speech segment and its corresponding auxiliary segment for the enrollment speech, while a different speaker is chosen for the interference speech. There is a 5% probability that no interfering speaker is present. Target speech and interfering speech are mixed with an SNR ranging from [-15, 15] dB, with an additional 10% probability of adding extra noise.
The audio length for each batch is 8 seconds. Before being fed into the model, the audio is randomly truncated to a length between 4 and 8 seconds to ensure the model’s ability to generalize to different audio lengths. During training, the distribution of tasks (NS, PLC, AEC, and TSE) is evenly balanced. Within each batch, the data are of the same task type. Gradient accumulation is enabled to help the model adapt to multi-task learning, with parameter updates occurring every 20 steps.
Model configuration. We use the open-source checkpoints of WavLM-large [^2] WavLM-Large on Hugging Face and X-codec2333X-codec2 on Hugging Face. The LLaMA-based LM comprises 16 LLaMA layers, each with 16 attention heads, a dropout rate of 0.1, a hidden size of 2048, and an intermediate size of 4096. The total number of parameters in the model is approximately 1.07 billion. More details are given in Appendix A.1.
Baseline systems. We evaluate the performance of our LLaSE-G1 with several state-of-the-art (SOTA) models of each subtask, including the winners of the recent signal processing grand challenges Reddy et al. (2020); Dubey et al. (2023); Cutler et al. (2023); Diener et al. (2024, 2022) for each task. Details of the baseline system and test set for each subtask are provided in Appendix A.
Evaluation Metrics. We use objective metrics to evaluate the performance of the baseline systems and our model. DNSMOS Reddy et al. (2022) include speech quality (SIG), background noise quality (BAK), and overall quality (OVRL) of the audio. AECMOS Purin et al. (2022) consists of echo annoyance MOS (EMOS) and other degradation MOS (DMOS). PLCMOS Diener et al. (2023) is used to assess the quality of audio processed by PLC algorithms. All MOS scores range from 1 to 5, representing audio quality from low to high. SpeechBERTScore (SBS) Saeki et al. (2024) is also employed to evaluate the semantic similarity between the enhanced audio and the reference audio. Following Zhang et al. (2025), we use HuBERT-base [^4] Hubert-base on Hugging Face model to extract semantic features. For acoustic similarity, we calculate speaker similarity SimWB based on the WavLM-base-sv model [^5] WavLM-base-sv on Huggingface to evaluate the performance.
Inference. For each task, we conduct single and multiple inferences. For multiple inferences, we infer 10 times and take the best result where the model’s output is used as the input for the next inference. For the PLC task, we only employ the audio to be processed as input, without the lossy label. For the TSE task, we keep the enrollment audio unchanged during multiple inferences. For the AEC task, we only use the reference audio for the first inference, subsequent inferences are treated as NS tasks. For the SS task, we employ a two-stage inference strategy. First, we separate one speaker from the mixed audio, and then use the first separated speaker’s audio as a reference for the second inference stage.
4.2 Experimental Results
4.2.1 Noise Suppression
Table 2 presents a comparison between the proposed LLaSE-G1 and several SOTA discriminative and generative models. The "With Reverb" column represents the test set with reverberation, and the "No Reverb" column is the one without. The results indicate that generative NS models consistently outperform discriminative models, especially in reverberant conditions. With single inference, LLaSE-G1 already surpassed most other systems. After multiple inferences, its performance improves further, achieving a SOTA result of 3.49 OVRL score at the no_reverb test set and 3.42 OVRL score at the with_reverb test set.
[…]
Model Type With Reverb No Reverb
SIG BAK OVRL SIG BAK OVRL
Noisy - 1.76 1.50 1.39 3.39 2.62 2.48
Conv-TasNet D 2.42 2.71 2.01 3.09 3.34 3.00
DEMUCS D 2.86 3.90 2.55 3.58 4.15 3.35
FRCRN D 2.93 2.92 2.28 3.58 4.13 3.34
SELM G 3.16 3.58 2.70 3.51 4.10 3.26
MaskSR G 3.53 4.07 3.25 3.59 4.12 3.34
AnyEnhance G 3.50 4.04 3.20 3.64 4.18 3.42
GenSE G 3.49 3.73 3.19 3.65 4.18 3.43
LLaSE-G1single G 3.59 4.10 3.33 3.66 4.17 3.42
LLaSE-G1multi G 3.65 4.16 3.42 3.71 4.19 3.49
4.2.2 Packet Loss Concealment
We compared LLaSE-G1 with the top-performing models Zhang et al. (2024); Li et al. (2022); Liu et al. (2022a); Valin et al. (2022) from the most recent two challenges on the Interspeech 2022 PLC blind test set Diener et al. (2022). It is important to note that LLaSE-G1 operates as a blind PLC without the need for lossy labels. This means LLaSE-G1 autonomously determines whether to perform PLC without prior knowledge of which frames experienced packet loss, making it a more challenging task and distinct from the models participating in the PLC Challenge.
Table 3: DNSMOS OVRL and PLCMOS scores on ICASSP 2022 PLC-challenge blind testet.
Model Type OVRL PLCMOS
Noisy - 2.56 2.90
KuaishouNet D - 4.27
LPCNet D 3.09 3.74
PLCNet D - 3.83
BS-PLCNet D 3.20 4.29
LLaSE-G1single G 3.03 3.68
LLaSE-G1multi G 3.27 4.30
The results in Table 3 demonstrate significant improvement with our model through inference time scaling. Specifically, the multi-inference approach boosts both OVRL and PLCMOS scores, with OVRL increasing from 3.03 to 3.27 and PLCMOS rising from 3.68 to 4.30, highlighting its effectiveness. LLaSE-G1’s results on blind PLC surpassed those of other models using informed PLC, demonstrating the powerful content understanding and generation capabilities of LMs.
4.2.3 Target Speaker Extraction
We use the ICASSP 2023 DNS blind test set Dubey et al. (2023) for the TSE task evaluation, which includes two tracks: the headset track and the speakerphone track.
Table 4: pDNSMOS scores on ICASSP 2023 DNS-challenge blind testet.
Model Type Track 1 Track 2
SIG BAK OVRL SIG BAK OVRL
Noisy - 4.15 2.37 2.71 4.05 2.16 2.50
TEA-PSE 3.0 D 4.12 4.05 3.65 3.99 3.95 3.49
NAPSE D 3.81 3.99 3.38 3.92 4.17 3.56
LLaSE-G1single G 4.21 3.99 3.72 4.08 3.84 3.55
LLaSE-G1multi G 4.20 3.97 3.70 4.11 3.86 3.58
As shown in Table 4, in both tracks, LLaSE-G1’s SIG MOS significantly outperformed other methods, indicating that LMs-based generative models provide higher audio quality with less signal distortion. The OVRL scores suggest that TEA-PSE 3.0 and NAPSE have advantages on headset and speakerphone devices, respectively. However, the proposed LLaSE-G1 achieved the best performance across both tracks, demonstrating superior device generalization capabilities compared to discriminative models.
4.2.4 Acoustic Echo Cancellation
LLaSE-G1 is the first generative model to integrate the AEC task into a unified framework. As shown in Table 5, LLaSE-G1 demonstrates comparable performance to the SOTA discriminative AEC approaches, showcasing the potential of LMs-based generative models for the AEC task.
Table 5: AECMOS Echo (EMOS) and Degradation (DMOS) scores on ICASSP 2023 AEC-challenge blind test set."DT" represents double-talk, FEST means far-end only and NEST means near-end only.
Model Type DT FEST NEST
EMOS DMOS EMOS DMOS
Align-CRUSE D 4.60 3.95 4.56 -
DeepVQE D 4.70 4.29 4.69 4.41
ULCNetAENR D 4.54 3.58 4.73 4.15
Align-ULCNet D 4.60 3.80 4.77 4.28
LLaSE-G1single G 4.42 3.82 4.64 3.66
LLaSE-G1multi G 4.52 3.91 4.65 3.50
(a) NS
(b) PLC
(c) AEC
(d) TSE
Figure 2: Inference-time scaling results on different tasks
4.2.5 Emergent Capabilities and Scaling Effects at Test Time
Emergent capabilities. The SS task is not included in the training data, we use it to test the emergent capabilities of LLaSE-G1. When compared to other discriminative methods, our LLaMA-based LLaSE-G1 demonstrates significant emergent capabilities. With multiple inferences, our generative model outperforms discriminative methods in OVRL scores of 3.17 and 3.25 on test sets, highlighting the potential of LLaSE-G1 to go beyond task-specific optimizations and adapt seamlessly to new tasks.
Table 6: DNSMOS scores on Libri2mix and WSJ0_2MIX test set.
Model Type Libri2mix WSJ0_2MIX
SIG BAK OVRL SIG BAK OVRL
Noisy - 2.33 1.66 1.64 3.42 3.20 2.76
Sepformer D 3.33 3.88 3.02 3.43 3.96 3.14
Mossformer2 D 3.44 3.94 3.11 3.50 4.05 3.23
LLaSE-G1single G 3.48 3.83 3.11 3.52 3.92 3.19
LLaSE-G1multi G 3.50 3.90 3.17 3.55 3.97 3.25
Inference-time scaling. As shown in Figure 2, scaling the inference time improves model performance across nearly all tasks. For the AEC and TSE tasks, performance peaks after the second inference, with EMOS improving from 4.42 to 4.52 and DMOS rising from 3.82 to 3.91. In contrast, the PLC task shows a significant performance boost with increased inference time, with PLCMOS rising from 3.67 to 4.30 and OVRL improving from 3.03 to 3.27, a gain of up to 25%. For the NS task, the OVRL score increases from 3.42 to 3.49 on the no_reverb dataset and from 3.33 to 3.42 on the with_reverb dataset. These results show that scaling test-time compute will initially improve performance, and decrease later due to the accumulation of acoustic distortion.
4.2.6 Semantic and Speaker Similarity
As shown in Table 7, we compare the semantic and speaker similarity between baseline systems and LLaSE-G1. Notably, TSE and AEC tasks are tested on the blind test sets where ground-truth speech is unavailable. So, we conduct evaluations of NS, PLC, and SS tasks. LLaSE-G1 outperforms generative SE models while getting slightly lower results in SBS, suggesting LLaSE-G1 effectively maintains speech content. Moreover, LLaSE-G1 achieves the highest SimWB in the NS task and competitive SimWB in the PLC and SS tasks, showing superior acoustic preservation capability.
Table 7: Semantic and speaker similarity results on various tasks, using the same test sets from previous subsections.
Task Model Type SBS SimWB
NS FRCRN D 0.85 0.980
AnyEnhance G 0.82 0.970
SELM G 0.72 0.965
GenSE G 0.78 0.974
LLaSE-G1 G 0.83 0.993
PLC BS-PLCNet D 0.95 0.999
LLaSE-G1 G 0.85 0.992
SS Sepformer D 0.85 0.980
Mossformer2 D 0.87 0.991
LLaSE-G1 G 0.82 0.988
4.2.7 Ablation Study
We conduct an ablation study to evaluate the effectiveness of input representations, output representations, and model backbone, choosing SELM as the baseline. As shown in Table 8, when replacing inputs and output with proposed continuous features and speech tokens, there is an obvious improvement, revealing the effectiveness of acoustic preservation. Besides, there is no performance drop when replacing the full attention Transformer with casual attention LLaMA. Finally, adopting a multi-inference strategy further boosts performance
Table 8: DNSMOS scores on DNS blind test set without reverb. "D" represents Discrete tokens, and "C" represents Continuous features. "S" represents Single inference, and "M" represents Multiple inference.
Input LM Output Inference OVRL
Noisy - - - - 2.48
Baseline D Transformer HiFiGAN S 3.26
C Transformer HiFiGAN S 3.34
C LLaMA HiFiGAN S 3.35
Proposed C LLaMA X-codec2 S 3.43
C LLaMA X-codec2 M 3.49
5 Conclusion
In this study, we propose LLaSE-G1, a general LLaMA-based framework to unify various speech enhancement tasks. Specifically, we employ continuous features as input and predict 1D speech tokens, maximizing acoustic preservation. Additionally, we design dual-channel inputs and outputs, unifying multiple SE tasks. Extensive experiments show that LLaSE-G1 achieves superior performance in each benchmark, serving as a powerful foundation model. Moreover, LLaSE-G1 demonstrates scaling effects at test time and emerging capabilities for unseen SE tasks.
Limitations
Although LLaSE-G1 demonstrates promising results across diverse SE tasks, there are several limitations that can be addressed towards LLaSE-G2. First, LLaSE-G1 operates at a 16,000 Hz sampling rate due to WavLM and X-codec2. We plan to support full-band audio and super-resolution generation in future research. Second, the training data and model size of LLaSE-G1 are relatively small as compared with that used in mainstream audio langauge models for understanding and conversation tasks. Hence we would like to further scale up data and model size to boost performance in generative speech enhancement.
Appendix A Appendix for Experimental Set Up
A.1 Model Configuration
We use the open-source checkpoints of WavLM-large and X-codec2. The LLaMA-based LM comprises 16 LLaMA layers, each with 16 attention heads, a dropout rate of 0.1, a hidden size of 2048, and an intermediate size of 4096. The total number of parameters in the model is approximately 1.07 billion, which includes all learnable weights and biases across all layers. The model has 2 input linear layers and 2 output linear layers. The input layer maps a 1024-dimensional vector to another 1024-dimensional vector, while the output layer transforms a 2048-dimensional vector into a 65536-dimensional vector, which is the codebook size of Xcodec2.
We trained the model for 100,000 steps using 4 NVIDIA L40 GPUs, with a batch size of 6 per GPU and the AdamW optimizer. The learning rate is set to 1e-4.
A.2 Test Sets
NS: Interspeech 2020 DNS Challenge blind Test Set. Reddy et al. (2020) It contains 600 clips (300 synthetic and 300 real), with synthetic clips generated using clean speech and noise not seen during training, and real clips crowdsourced in diverse noisy conditions.
PLC: Interspeech 2022 PLC Challenge test set Diener et al. (2022) This is a realistic evaluation dataset based on packet loss patterns from actual calls, providing a methodology for comparing different approaches and a new objective metric to help researchers improve their techniques.
TSE: ICASSP 2023 DNS Challenge blind Test Set Dubey et al. (2023) The blind test set includes two tracks Headset and Speakerphone with clips featuring 10-30 seconds of enrollment speech, with or without noise. It is used for final rankings and evaluates both personalized and non-personalized models using the Personalized ITU-T P.835 framework.
AEC: ICASSP 2023 AEC Challenge blind Test Set Cutler et al. (2023) The blind test set in the AEC Challenge consists of real-world data collected from over 10,000 diverse audio devices and environments. It is used to determine the final competition winners. The dataset includes recordings of both single-talk and double-talk scenarios, with varying conditions like background noise, reverberation, and device distortions.
SS: Libri2mix Cosentino et al. (2020), WSJ0-2mix. These two test sets are commonly used in speech separation, which is mixed from librispeech and WSJ datasets.
A.3 Baseline Systems
NS: For discriminative systems, we choose Conv-TasNet Luo and Mesgarani (2019),DEMUCS  Défossez et al. (2019),FRCRN Zhao et al. (2024a) , which is recent SOTA models on noise suppression. For generative systems, we choose SELM Wang et al. (2024), which introduce LM to speech enhancement, and GenSE Yao et al. (2025) and AnyEnhance Zhang et al. (2025), 2 newly released SOTA-level generative speech enhancement systems.
PLC: we use BS-PLCNet  Zhang et al. (2024), Team Kuaishow Li et al. (2022), which are the winners of the 2024 challenge and 2022 challenge respectively, and other systems in challenge like PLCNet Liu et al. (2022a) and LPCNet Valin et al. (2022) as our baseline systems.
TSE: We compare our model with two baseline systems: TEA-PSE 3.0 Ju et al. (2023), the winner of the challenge, and NAPSE Yan et al. (2023), which placed second.
AEC: For baseline comparison, we choose recent efficient and state-of-the-art systems as our baseline systems, including UCLNet, Shetu et al. (2024a),AlignUCLNet Shetu et al. (2024b),AlignCruse Indenbom et al. (2023b) and DeepVQE Indenbom et al. (2023a), which is the state-of-the-art model in the AEC task.
SS: We use SOTA discriminative speech separation systems such as Sepformer Subakan et al. (2021) and Mossformer2 Zhao et al. (2024c), which is the SOTA system on speech separation, as our SS baseline systems.